Please help me solve this problem: If you buy a new TV for $599.25 on sale, which is 25% off of the regular price, and the store marked the price up by 100%, what was the store's cost for the TV?

Answer by stanbon(75887)

Please help me solve this problem: If you buy a new TV for $599.25 on sale, which is 25% off of the regular price, and the store marked the price up by 100%, what was the store's cost for the TV?
-------------
Let the store's cost be "x".
Price after markup = "2x"
---
Discounted price = 0.75*2x = 1.5x
------
Equation:
1.5x = 599.25
---
x = 599.25/1.5 = $399.50 (store's cost)
-------------------------------------------
Cheers,
Stan H.
---------------